# Irreducibility and periodicity in $\mathbb{Z}^{2}$ symbolic systems

### v1

@@ 5.2 Proof setup
By our assumptions, $\widehat{E}$ separates $E_{x}$ from $E_{y}$ and is at least $1$-far from each of them. We call $\widehat{E}$ the gap. We call $\widehat{E}_{x}$ the zone of $C_{x}$ and $\widehat{E}_{y}$ the zone of $E_{y}$. For a frame $F$ in one of the certificates $C_{x},C_{y}$, we write $C_{F}=C_{x}$ or $C_{y}$ depending on whether $F\in C_{x}$ to $F\in C_{y},$and similarly $E_{F}=E_{x}$ or $E_{y}$ and $\widehat{E}_{F}=\widehat{E}_{x}$ or $\widehat{E}_{y}$. We write $C^{\prime}_{F}$ for the certificate that $F$ does not belong to and similarly $E^{\prime}_{F}\widehat{E}^{\prime}_{F}$. We use similar notation for boxes and witnesses.
@@
Both the extension of $z$ and the definition of $C$ will be carried out in an iterative fashion. We begin from $C=\emptyset$, which is trivially compatible with $z$. Frames will then be added to $C$ one at a time in several (infinite) rounds. Initially, we add frames from $C_{x}$ or $C_{y}$ that require little or no modification in order to maintain the certificate properties and be compatible with $z$. Afterwards, we add frames from $C_{x},C_{y}$ that require more extensive changes, and finally we add entirely new frames, although they will also be derived, after substantial changes, from frames in $C_{x}C_{y}$.
@@ 5.3 Performing the gluing
      If $\gamma\cap\widehat{E}=\emptyset$, we set $w^{\prime}=w$ (i.e. we do not “move” $w$ at all).
@@
Having added $w^{\prime}$ to each box $B\in F^{\prime}$ we claim now that $C\cup\{F^{\prime}\}$ is a certificate. Since the boxes in $F^{\prime}$ come from those of $F$ they satisfy all the relevant conditions from the definition. It remains to check two things:
  • The center of $F^{\prime}$ (equivalently, of $F$) is $r_{n}/2$-far from the centers of other $n$-frames $G$ in $C$
@@
    Consider a witness $w^{\prime}\in F^{\prime}$ belonging to the section $\sigma=\sigma(w^{\prime})$ of an $n$-box $B^{\prime}\in F^{\prime}$. :et $w,\gamma$ be as defined above when $w^{\prime}$ was added.
@@
    Thus, by part (1) of 3.1, it suffices to show that for $k<n$, if $S\in\mathcal{R}_{C\cup\{F^{\prime}\}}(B^{\prime})\setminus\mathcal{R}_{C_{F}}(B^{\prime})$, then $w^{\prime}\notin 20S$.
    Suppose to the contrary there exists a level-$k$ box $B^{\prime\prime}\in C\cup\{F^{\prime}\}$ belonging to a frame $F^{\prime\prime}\in C\cup\{F^{\prime}\}$, and a section $S=(B^{\prime\prime})^{(i)}$, such that $S\in\mathcal{R}_{C\cup F}\setminus\mathcal{R}_{C_{F}}$ and $w^{\prime}\in 20S$.
@@
    • Since $F^{\prime\prime}$ was added to $C$ in Step A, it intersects $\widehat{E}_{F^{\prime\prime}}$ (its own zone, which is different from that of $F$).
@@
    Thus, we can travel along $\gamma$ from $w^{\prime}$ towards $c(F)$, we will pass within $\frac{1}{100}r_{k}+20h_{k}+\frac{1}{1000}r_{k}$ of the center of $F^{\prime\prime}$, and this distance can be assumed smaller than $\frac{1}{10}r_{k}$. But every point within $\frac{1}{10}r_{k}$ of the center of $F^{\prime\prime}$ lies in the zone $\widehat{E}_{F^{\prime\prime}}$, because if $F^{\prime\prime}$ was added to $C$ in stage A. This proves the claim. ∎
@@
We must verify that $C\cup\{F^{\prime}\}$ is a certificate. We havea added frames at different levels than step A and all frames and boxes we added come from $C_{x}$, so separation of boxes and frames is inheritted from $C_{x}$. We also can not have disrupted the safety property of existing witnesses from Step A because we are adding only higher levels than before. What we must check is that the new witnesses added in the present stage satisfy the safety property.
@@
Proof.
Proof of the Claim] By our choice of $w^{\prime}$, if $w^{\prime}\notin\widehat{E}$ then $\gamma$ does not intersect $\widehat{E}$ at all. So, if $w^{\prime}\in\widehat{E}_{y}$ then $w^{\prime}\notin\widehat{E}$, and, since $\widehat{E}$ separates $\widehat{E}_{x},\widehat{E}_{y}$, we must also have $\gamma\cap\widehat{E}_{x}=\emptyset$. It follows that if $w^{\prime}\in\widehat{E}_{y}$, then $\gamma\subseteq\widehat{E}_{y}$. On the other hand $w^{\prime}$ is an $n$-witness for $n\geq n_{0}$, so the diameter of $\widehat{E}$ (and hence of $\widehat{E}\cup\widehat{E}_{y}$) is less than $\frac{1}{10}r_{n}$, while the diameter of $\gamma$ is at least $w_{n}=\frac{9}{10}r_{n}$, so $\gamma\subseteq\widehat{E}_{y}$ is impossible, and consequently so is $w^{\prime}\in\widehat{E}_{y}$. ∎
@@
    But $\gamma^{\prime}\cap\widehat{E}=\emptyset$, since any point in $\gamma^{\prime}\cap\widehat{E}$ would be closer to $c(F)$ than $w^{\prime}$, contradicting the definition of $w^{\prime}$.
@@
    Since $n\geq n_{0}$, it follows that $\widehat{E}_{y}$, and hence also $\gamma^{\prime}$, has diameter $<r_{n}/10$.
    Therefore the distance of $w^{\prime}$ (one end of $\gamma^{\prime}$) from the short edge of $B$ that is closer to $c(F)$, is $<\frac{1}{10}r_{n}=\frac{1}{9}w_{n}$.
@@
Both cases lead to a contradiction, so $w^{\prime}\notin 20S$ and hence $w^{\prime}$ is safe relative to $\mathcal{R}_{C\cup\{F^{\prime}\}}(B)$. We have shown that $C\cup\{F^{\prime}\}$ is a certificate.
@@
First, we observe that adding frames with these centers does not violate the $r_{n}/2$ separation condition of $n$-frames. Indeed, since $n\leq n_{0}$, at the end of Step C all $n$-frames in $C$ have come from steps A and C of the construction (since Step B only adds frames at levels higher than $n_{0}).$ Frames that cam from the same certificate are separated by assumption, whereas if they came from different certificates then the centers are in their own zones and $r_{n}/4$-far from $\widehat{E}$, hence they are $r_{n}/2$ from each other (see Step A item (1)).
@@
Then $W_{>n}$ contains at most one witness $w$, because $h_{n+1}\gg r_{n}$, and by Lemma 4.2, the witnesses at levels $>n$ are $h_{n+1}$-separated, while all the rectangles $R_{i}$ (and hence their sections) lie within a single $r_{n}$-ball.
Because the $R_{i}$ are $100h_{n}$-separated, it follows that at most for all but at most one of them, all sections $S$ of $R_{i}$ satisfy $w\notin 20S$. If there is an exceptional one, throw it out.
@@
Taking stock, for $n\geq n_{0}$ the centers of $n$-frames in $C$ are the same as the centers of $n$-frames in $C_{x}$, so they are $10r_{n}$ dense. But for $n\leq n_{0}$, although we have added to $C$ many frames from $C_{x},C_{y}$, we have not added those whose centers come too close to the opposite zones. Therefore, for $1\leq n<n_{0}$, the centers of $n$-frames in $C$ may not be $10r_{n}$-dense.
We cannot directly add the missing frames from levels $n\leq n_{0}$ because, being close to the opposite zone, they may come to close to frames from the opposite zone that are alreay in $C$.
@@
Let $\ell$ denote the ray starting at $u$ and passing through $v$, and let $u^{\prime}\in\ell$ be the point closest to $u$ that satisfies $d(e,u^{\prime})=\frac{3}{4}r_{n}$. Note that $d(u^{\prime},u)\geq\frac{1}{2}r_{n}$ because after moving a distance $t<\frac{1}{2}r_{n}$ along $\ell,$we are at a point that is less than $\frac{1}{4}r_{n}+t\leq\frac{1}{4}r_{n}+\frac{1}{2}r_{n}<\frac{3}{4}r_{n}$ from $e$.
@@ 6 Concluding remarks and problems
Indeed, one can show that our example contains points with one direction of periodicity, so it does not provide an answer to the problem above.
@@
where $N_{n}(x)$ is the number of $n\times n$ patterns in $x$
The existence of a sysmem satisfying the first condition with $\alpha>1$ would give a positive solution to the previous problem. At the other extreme, a system that fails the second condition for $\alpha=2$ and small enough $c$ must contain periodic points, by results related to Nivat’s conjecture.
@@
Problem 6.3.
Let $d\geq 3$. If $X$ is a SI $\mathbb{Z}^{d}$ SFT and $Z$ is an infinite free minimal $\mathbb{Z}^{d}$-subshift and $h_{top}(Z)<h_{top}(X)$, is there an embedding (an injective factor map) $Z\rightarrow X$

### v2

@@ 6.2 Proof setup
By our assumptions, $\widehat{E}$ separates $E_{x}$ from $E_{y}$ and is at least $1$-far from each of them. We call $\widehat{E}$ the gap. We call $\widehat{E}_{x}$ the zone of $E_{x}$ and $\widehat{E}_{y}$ the zone of $E_{y}$. For a frame $F$ in one of the certificates $C_{x},C_{y}$, we write $C_{F}=C_{x}$ or $C_{y}$ depending on whether $F\in C_{x}$ to $F\in C_{y}$, and similarly $E_{F}=E_{x}$ or $E_{y}$ and $\widehat{E}_{F}=\widehat{E}_{x}$ or $\widehat{E}_{y}$. We write $C^{\prime}_{F}$ for the certificate that $F$ does not belong to and similarly $E^{\prime}_{F},\widehat{E}^{\prime}_{F}$. We use similar notation for boxes and witnesses.
@@
Both the extension of $z$ and the definition of $C$ will be carried out in an iterative fashion. We begin from $C=\emptyset$, which is trivially compatible with $z$. Frames will then be added to $C$ one at a time in several (infinite) rounds. Initially, we add frames from $C_{x}$ or $C_{y}$ that require little or no modification in order to maintain the certificate properties and be compatible with $z$. Afterwards, we add frames from $C_{x},C_{y}$ that require more extensive changes, and finally we add entirely new frames, although they will also be derived, after substantial changes, from frames in $C_{x},C_{y}$.
@@ 6.3 Performing the gluing
      If $\gamma\cap\widehat{E}$$\emptyset$, we set $w^{\prime}=w$ (i.e. we do not “move” $w$ at all).
@@
Having added a $w^{\prime}$ to each box $B\in F^{\prime}$ we claim now that $C\cup\{F^{\prime}\}$ is a certificate. Since the boxes in $F^{\prime}$ come from those of $F$ they satisfy all the relevant conditions from the definition. It remains to check two things:
  • The center of $F^{\prime}$ (equivalently, of $F$) is $r_{n}/2$-far from the centers of other $n$-frames $G$ in $C$. There are two cases:
@@
    Consider a witness $w^{\prime}\in F^{\prime}$ belonging to the section $\sigma=\sigma(w^{\prime})$ of an $n$-box $B^{\prime}\in F^{\prime}$. Let $w,\gamma$ be as defined above when $w^{\prime}$ was added.
@@
    Thus, by part (1) of Proposition 4.1, it suffices to show that for $k\leq n$, if $S\in\mathcal{R}_{C\cup\{F^{\prime}\}}(B^{\prime})\setminus\mathcal{R}_{C_{F}}(B^{\prime})$, then $w^{\prime}\notin 20S$.
    The case $k=n$ follows directly from Lemma 5.1, because the fact that $F^{\prime},F^{\prime\prime}$ come from different certificates and their centers belong to the corresponding zones and are at least $\frac{2}{3}r_{n}$ far from the complementary zone, means that their centers are more than $\frac{1}{2}r_{n}$ apart, as required by the lemma.
    Now suppose that $1\leq k<n$ and that there exists a level-$k$ box $B^{\prime\prime}\in C\cup\{F^{\prime}\}$ belonging to a frame $F^{\prime\prime}\in C\cup\{F^{\prime}\}$, and a section $S=(B^{\prime\prime})^{(i)}$, such that $S\in\mathcal{R}_{C\cup\{F^{\prime}\}}(B^{\prime})\setminus\mathcal{R}_{C_{F}}(B^{\prime})$ and $w^{\prime}\in 20S$.
@@
    • Since $F^{\prime\prime}$ was added to $C$ in Step A, it intersects $\widehat{E}_{F^{\prime\prime}}$ (its own zone, which is different from that of $F^{\prime}$).
@@
    Thus, we can travel along $\gamma$ from $w^{\prime}$ towards $c(F)$, we will pass within $\frac{1}{100}r_{k}+20h_{k}+\frac{1}{1000}r_{k}$ of the center of $F^{\prime\prime}$, and this distance can be assumed smaller than $\frac{1}{10}r_{k}$. But every point within $\frac{1}{10}r_{k}$ of the center of $F^{\prime\prime}$ lies in the zone $\widehat{E}_{F^{\prime\prime}}$, because if $F^{\prime\prime}$ was added to $C$ in step A. This proves the claim. ∎
@@
We must verify that $C\cup\{F^{\prime}\}$ is a certificate. We have added frames at different levels than step A and all frames and boxes we added come from $C_{x}$, so separation of boxes and frames is inherited from $C_{x}$. We also can not have disrupted the safety property of existing witnesses from Step A because we are adding only higher levels than before. What we must check is that the new witnesses added in the present stage satisfy the safety property.
@@
Proof of the Claim.
By our choice of $w^{\prime}$, if $w^{\prime}\notin\widehat{E}$ then $\gamma$ does not intersect $\widehat{E}$ at all. So, if $w^{\prime}\in\widehat{E}_{y}$ then $w^{\prime}\notin\widehat{E}$, and, since $\widehat{E}$ separates $\widehat{E}_{x},\widehat{E}_{y}$, we must also have $\gamma\cap\widehat{E}_{x}=\emptyset$. It follows that if $w^{\prime}\in\widehat{E}_{y}$, then $\gamma\subseteq\widehat{E}_{y}$. On the other hand $w^{\prime}$ is an $n$-witness for $n\geq n_{0}$, so the diameter of $\widehat{E}$ (and hence of $\widehat{E}\cup\widehat{E}_{y}$) is less than $\frac{1}{10}r_{n}$, while the diameter of $\gamma$ is at least $w_{n}=\frac{9}{10}r_{n}$, so $\gamma\subseteq\widehat{E}_{y}$ is impossible, and consequently so is $w^{\prime}\in\widehat{E}_{y}$. ∎
@@
    But $\gamma^{\prime}\cap\widehat{E}$$\emptyset$, since any point in $\gamma^{\prime}\cap\widehat{E}$ would be closer to $c(F)$ than $w^{\prime}$, contradicting the definition of $w^{\prime}$.
@@
    Since $n\geq n_{0}$, it follows that $\widehat{E}_{y}$, and hence also $\gamma^{\prime}$, has diameter $r_{n}/10$.
    Therefore the distance of $w^{\prime}$ (one end of $\gamma^{\prime}$) from the short edge of $B$ that is closer to $c(F)$, is $\frac{1}{10}r_{n}=\frac{1}{9}w_{n}$.
@@
dvBoth cases lead to a contradiction, so $w^{\prime}\notin 20S$ and hence $w^{\prime}$ is safe relative to $\mathcal{R}_{C\cup\{F^{\prime}\}}(B)$. We have shown that $C\cup\{F^{\prime}\}$ is a certificate.
@@
First, we observe that adding frames with these centers does not violate the $r_{n}/2$ separation condition of $n$-frames. Indeed, since $n\leq n_{0}$, at the end of Step C all $n$-frames in $C$ have come from steps A and C of the construction (since Step B only adds frames at levels higher than or equal to $n_{0}).$ Frames that came from the same certificate are separated by assumption, whereas if they came from different certificates then the centers are in their own zones and $r_{n}/4$-far from $\widehat{E}$, hence they are $r_{n}/2$ from each other (see Step A item (1)).
@@
Then $W_{>n}$ contains at most one witness $w$, because $h_{n+1}\gg r_{n}$, and by Lemma 5.2, the witnesses at levels >n>n are $h_{n+1}$-separated, while all the rectangles $R_{i}$ (and hence their sections) lie within a single $r_{n}$-ball.
Because the $R_{i}$ are $100h_{n}$-separated, it follows that for all but at most one of them, all sections $S$ of $R_{i}$ satisfy $w\notin 20S$. If there is an exceptional one, throw it out.
@@
Taking stock, for $n\geq n_{0}$ the centers of $n$-frames in $C$ are the same as the centers of $n$-frames in $C_{x}$, so they are $10r_{n}$ dense. But for $n<n_{0}$, although we have added to $C$ many frames from $C_{x},C_{y}$, we have not added those whose centers come too close to the opposite zones. Therefore, for $1\leq n<n_{0}$, the centers of $n$-frames in $C$ may not be $10r_{n}$-dense.
We cannot directly add the missing frames from levels $\leq n_{0}$ because, being close to the opposite zone, they may come too close to frames from the opposite zone that are already in $C$.
@@
[…]
@@ 7 Concluding remarks and problems
Indeed, one can show that our example contains points with one direction of periodicity, so it does not provide an answer to the problem above. Ronnie Pavlov [12] has an example of a free square mixing system, but this is still far from strong irreducibility.
@@
where $N_{n}(x)$ is the number of $n\times n$ patterns in $x$ ?
The existence of a system satisfying the first condition with $\alpha>1$ would give a positive solution to the previous problem. At the other extreme, a system that fails the second condition for $\alpha=2$ and small enough $c$ must contain periodic points, by results related to Nivat’s conjecture.
@@
Problem 7.3.
Let $d\geq 3$. If $X$ is a SI $\mathbb{Z}^{d}$ SFT and $Z$ is an infinite free minimal $\mathbb{Z}^{d}$-subshift and $h_{top}(Z)<h_{top}(X)$, is there an embedding (an injective factor map) $Z\rightarrow X$ ?